OpenStudy (anonymous):

MEDAL What would 10/25 be A) 3/4 B) 2:5 C) 65% D) 6 to 16

OpenStudy (unklerhaukus):

do you know factors of 10?

OpenStudy (anonymous):

um 5, 10

OpenStudy (anonymous):

1, 2, 5, 10

OpenStudy (unklerhaukus):

that's right

OpenStudy (anonymous):

its 2:5

OpenStudy (unklerhaukus):

what about factors of 25?

OpenStudy (anonymous):

1,2,3,4,9,12,18,36

OpenStudy (unklerhaukus):

?

OpenStudy (anonymous):

is it 2:5

OpenStudy (unklerhaukus):

the factors of 25 are 1, 5 , 5 and 25

OpenStudy (unklerhaukus):

So we have established that 10 = 2*5 and that 25 = 5*5 Lets apply this to the fraction 10/25 = (2*5) / (5*5) now we can cancel the common factor = 2 / 5 which is as you said equivalent to the ratio 2:5